# Versor: A Geometric Sequence Architecture Enhanced Scale Generalization and Interpretability via Conformal Algebra

### v1

@@ B.4.2 Riemannian Gradient Correction
By projecting the state $\Psi$ onto the unit hypersphere at every step, we approximate Riemannian optimization by ensuring the point remains on the manifold. This ensures that Manifold Normalization is not just a numerical stabilizer, but a computationally cheap proxy for Riemannian optimization.
@@ C.1 Derivation of the Bit-Masked Geometric Product
We define an isomorphism $\phi:\mathcal{G}\to\mathbb{Z}_{2}^{n}$ between the Grassmann basis $\mathcal{G}$ and the $n$-dimensional bit-field. For any blade $e_{S}$, $\phi(e_{S})=\sum_{k\in S}2^{k}$. The geometric product $e_{i}e_{j}$ is then mapped to the bitwise domain as:
@@
Let the basis vectors $\{e_{1},e_{2},e_{3},e_{+},e_{-}\}$ be mapped to indices $\{0,1,2,3,4\}$. We represent any basis blade $e_{S}$ describing a subspace $S$ as an integer bitmask $i\in\{0,\dots,31\}$, where:
@@
The geometric product of two blades is defined by the juxtaposition of their basis vectors, $e_{i}e_{j}$. We assume canonical ordering where basis vectors in defined blades are strictly unique and ordered (e.g., $e_{1}e_{2}$ not $e_{2}e_{1}$). Vectors appearing in both blades contract (squares to scalars), while unique vectors remain. This is equivalent to the Symmetric Difference of the sets of basis indices. In bitmask space, the XOR operator ($\oplus$) functionally implements the symmetric difference:
@@
This proves that the resulting blade’s basis is always found at the XORed index, eliminating the need for a search or hash map. Note that we additionally require the null basis relations $e_{o}\cdot e_{\infty}=-1$ which are handled in the metric signature weight.
@@
    To calculate $e_{i}e_{j}$, we move $e_{b_{1}}$ past all $e_{a_{k}}$ where $\text{index}(a_{k})>\text{index}(b_{1})$. Each such move incurs a sign change (1)(-1) due to $e_{a}e_{b}=-e_{b}e_{a}$.
@@
Bitwise Optimization: For a fixed bit $b\in j$, the number of bits in $i$ that are “to the left” (higher index) of $b$ is given by popcount(i & ~( (1 << (b+1)) - 1 )). Summing this over all bits in $j$ provides the total swap parity. A more efficient parallel version used in our Triton kernel is:
@@
    For each bit $k$ set in $mask_{intersect}$, we multiply the result by the signature $\eta_{kk}$.
@@ C.2 Computational Efficiency of the Bit-Masked Kernel
The Bit-Masked Method ($T_{bit}$): We iterate through all pairs of non-zero coefficients. In the dense case, there are $D^{2}$ pairs. For each pair, we compute the sign and index using bit-logic.
@@
For CGA ($32/5$), our method reduces the raw number of operations by a factor of 6.4x.
@@ C.4 The Memory Wall: Latency Proof
$\beta=\frac{Latency_{Table\_Lookup}}{n\cdot Latency_{ALU}}\approx\frac{40\text{ cycles}}{5\cdot 1\text{ cycle}}\approx\mathbf{8x}$
@@ C.6 Final Cumulative Speedup Calculation
$S_{total}\approx 6.4\times 8\approx\mathbf{51.2x}$
C.7 Proof of Isometric Conformal Embedding
@@
Theorem (Eq. 2): The inner product of two null vectors $X_{1},X_{2}\in\mathcal{C}l_{4,1}$ corresponds to the negative half-squared Euclidean distance between their original points $x_{1},x_{2}\in\mathbb{R}^{3}$.
@@
    Recall the Lifting $\mathcal{K}(\mathbf{x})$ (Eq. 1): $X=x+\frac{1}{2}x^{2}e_{\infty}+e_{o}$, where $e_{\infty}^{2}=0,e_{o}^{2}=0,$ and $e_{\infty}\cdot e_{o}=-1$.
@@
This verifies that Versor can calculate distances via linear dot products without ever using a $\sqrt{\cdot}$ or non-linear activation, explaining its efficiency in physics tasks.
C.8 Algorithmic Implementation Specifications
@@ 2 $B_{t}\leftarrow\text{ProjectToBivector}(u_{t},W_{B})$;
C.8.1 Theoretical Roofline Analysis
@@
C.8.2 Arithmetic Intensity ($\mathcal{I}$)
@@
Scenario B: Sparse Table Lookup (Naive GA) Standard GA libraries (Cliffod, PyGAE) use a sparse multiplication table.
@@
Scenario C: Versor Bit-Masked Kernel Our kernel computes indices/signs in registers.
@@
By converting Memory operations (Table Lookups) into ALU operations (Bitwise Logic), the algorithm is moved from the memory-bound region to the compute-bound region. On modern GPUs (P100, M4), ALUs are 100x faster than Memory, resulting in the substantial speedups observed.
@@ D.1 Derivation of Versor Initialization
Problem Formulation: In a standard neural network, weights are initialized to preserve the variance of activations across layers [13]. For a linear layer $y=Wx$, this requires $\text{Var}(y)\approx\text{Var}(x)$. We generalize this to the geometric product $y=Wx$ in $\mathcal{C}l_{4,1}$ (dimension $D=32$). Let $x=\sum_{i=1}^{D}x_{i}e_{i}$ and $W=\sum_{j=1}^{D}w_{j}e_{j}$, where $x_{i},w_{j}$ are independent and identically distributed random variables with mean 0 and variances $\sigma_{x}^{2},\sigma_{w}^{2}$. Recall that $\text{Var}(x)=E[x^{2}]-(E[x])^{2}=E[x^{2}]$. The geometric product is:
@@ D.2 Geometric Product Attention (GPA) Decomposition
    Scalar Part ($\langle\cdot\rangle_{0}$): From Appendix C.7, we know $Q\cdot K=-\frac{1}{2}d^{2}$ (approx). Thus, $\text{softmax}(\langle Q\widetilde{K}\rangle_{0})$ recovers the standard RBF-like distance attention used in Vanilla Transformers.
@@
    Torque Modulation: In our implementation, we incorporate the bivector magnitude $\|\langle Q\widetilde{K}\rangle_{2}\|$ directly into the attention score (see Eq. 4). This allows the model to prioritize interactions not just by proximity (scalar) but by the orientation and “torque” of the plane spanned by the interacting entities. Unlike Euclidean transformers which are blind to relative orientation, GPA attends to the full geometric configuration.
@@ E.3 Gradient Norm Preservation (Backward Stability)
During Backpropagation through Time (BPTT), we compute the gradient of the loss $\mathcal{L}$ with respect to the state $\Psi_{t}$ via the chain rule:
@@
The Jacobian matrix $J_{t}=\frac{\partial\Psi_{t+1}}{\partial\Psi_{t}}$ represents the linear transformation applied to $\Psi_{t}$. In our architecture, this transformation is the left-multiplication by the rotor $\Delta R_{t}$.
@@ E.4 Manifold Adherence of the Cayley Transform
Derivation: Let $B$ be a bivector. In $\mathcal{C}l_{4,1}$, the reversion of a bivector is $\widetilde{B}=-B$. We wish to show that $R\widetilde{R}=1$.
@@ Appendix F Experimental Setup and Detailed Results
The library abstracts the complexity of the bit-masked kernels behind a Pythonic API, an example of its use is below:
@@ F.1 Chaotic Dynamics (N-Body)
Extended Hamiltonian Details: The Hamiltonian-Versor Hybrid parameterizes the scalar energy function $H(q,p)$ using the Versor backbone. Input state $(q,p)\in\mathbb{R}^{6N}$ is lifted to $\mathcal{C}l_{4,1}$. The output scalar $\langle\Psi_{out}\rangle_{0}$ is treated as the Hamiltonian $\mathcal{H}$. Integration uses a hand-coded custom 4th-order Runge-Kutta symplectic integrator. This ensures $\frac{d\mathcal{H}}{dt}=0$ by construction (within integrator error).

### v2

@@ B.4.2 Riemannian Gradient Correction
By projecting the state $\Psi$ onto the unit hypersphere at every step, Riemannian optimization is approximated by ensuring the point remains on the manifold. This ensures that Manifold Normalization is not just a numerical stabilizer, but a computationally cheap proxy for Riemannian optimization.
@@ C.1 Derivation of the Bit-Masked Geometric Product
An isomorphism $\phi:\mathcal{G}\to\mathbb{Z}_{2}^{n}$ is defined between the Grassmann basis $\mathcal{G}$ and the $n$-dimensional bit-field. For any blade $e_{S}$, $\phi(e_{S})=\sum_{k\in S}2^{k}$. The geometric product $e_{i}e_{j}$ is then mapped to the bitwise domain as:
@@
Let the basis vectors $\{e_{1},e_{2},e_{3},e_{+},e_{-}\}$ be mapped to indices $\{0,1,2,3,4\}$. Any basis blade $e_{S}$ describing a subspace $S$ is represented as an integer bitmask $i\in\{0,\dots,31\}$, where:
@@
The geometric product of two blades is defined by the juxtaposition of their basis vectors, $e_{i}e_{j}$. Canonical ordering is assumed where basis vectors in defined blades are strictly unique and ordered (e.g., $e_{1}e_{2}$ not $e_{2}e_{1}$). Vectors appearing in both blades contract (squares to scalars), while unique vectors remain. This is equivalent to the Symmetric Difference of the sets of basis indices. In bitmask space, the XOR operator ($\oplus$) functionally implements the symmetric difference:
@@
This proves that the resulting blade’s basis is always found at the XORed index, eliminating the need for a search or hash map. Note that the null basis relations $e_{o}\cdot e_{\infty}=-1$ are additionally required and handled in the metric signature weight.
@@
    To calculate $e_{i}e_{j}$, $e_{b_{1}}$ is moved past all $e_{a_{k}}$ where $\text{index}(a_{k})>\text{index}(b_{1})$. Each such move incurs a sign change (1)(-1) due to $e_{a}e_{b}=-e_{b}e_{a}$.
@@
Bitwise Optimization: For a fixed bit $b\in j$, the number of bits in $i$ that are “to the left” (higher index) of $b$ is given by popcount(i & ~( (1 << (b+1)) - 1 )). Summing this over all bits in $j$ provides the total swap parity. A more efficient parallel version used in the Triton kernel is:
@@
    For each bit $k$ set in $mask_{intersect}$, the result is multiplied by the signature $\eta_{kk}$.
@@ C.2 Computational Efficiency of the Bit-Masked Kernel
The Bit-Masked Method ($T_{bit}$): All pairs of non-zero coefficients are iterated through. In the dense case, there are $D^{2}$ pairs. For each pair, the sign and index are computed using bit-logic.
@@
For CGA ($32/5$), the proposed method reduces the raw number of operations by a factor of 6.4x.
@@ C.4 The Memory Wall: Latency Proof
$\beta=\frac{Latency_{Table\_Lookup}}{n\cdot Latency_{ALU}}\approx\frac{60\text{ cycles}}{5\cdot 1\text{ cycle}}\approx\mathbf{12x}$
@@ C.6 Final Cumulative Speedup Calculation
$S_{total}\approx 6.4\times 12\approx\mathbf{76.8x}$
(Typically observed as $\sim 78\times$ in empirical benchmarks).
C.7 Matrix Isomorphism Acceleration ($\mathcal{C}l_{4,1}$)
To achieve the 1.05 ms per-step latency result, the periodicity of Clifford algebras is exploited. The $\mathcal{C}l_{4,1}$ algebra is isomorphic to a $4\times 4$ complex matrix algebra:
$\mathcal{C}l_{4,1}\cong\text{Mat}(4,\mathbb{C})$ (31)
This isomorphism allows for mapping the 32-dimensional multivector space of $\mathcal{C}l_{4,1}$ onto the 32-dimensional real space of $4\times 4$ complex matrices (since $4^{2}\times 2=32$).
C.7.1 Formal Derivation of the Representation
To construct the mapping, the representation of the generators $\{e_{i}\}_{i=1}^{5}$ is defined using Pauli matrices $\sigma_{x},\sigma_{y},\sigma_{z}$ and the $2\times 2$ identity $I_{2}$. The Kronecker product $\otimes$ is employed to build the $4\times 4$ matrices:
$\rho(e_{1})$ $=\sigma_{x}\otimes I_{2}$ (32)
$\rho(e_{2})$ $=\sigma_{y}\otimes I_{2}$ (33)
$\rho(e_{3})$ $=\sigma_{z}\otimes\sigma_{x}$ (34)
$\rho(e_{4})$ $=\sigma_{z}\otimes\sigma_{y}$ (35)
$\rho(e_{5})$ $=i(\sigma_{z}\otimes\sigma_{z})$ (36)
It is trivial to verify that these matrices satisfy the Clifford anticommutation relation $\{\rho(e_{i}),\rho(e_{j})\}=2\eta_{ij}I_{4}$, where $\eta=\text{diag}(+1,+1,+1,+1,-1)$. Specifically, $\rho(e_{1})^{2}=\dots=\rho(e_{4})^{2}=I_{4}$ and $\rho(e_{5})^{2}=-I_{4}$.
Any multivector $\Psi\in\mathcal{C}l_{4,1}$ can be written as a sum over the basis blades $E_{J}$:
$\Psi=\sum_{J\in\mathcal{P}\{1,\dots,5\}}\psi_{J}E_{J}\implies\rho(\Psi)=\sum_{J}\psi_{J}\rho(E_{J})$ (37)
where $\rho(E_{J})$ is the product of the corresponding generator matrices.
Computational Advantage. A naive geometric product in $\mathcal{C}l_{4,1}$ involves 1,024 floating-point operations (FLOPs) using the Cayley table. However, since the geometric product is equivalent to standard matrix multiplication in the representation space, the cost of multiplying two $4\times 4$ complex matrices is:
$4^{3}\text{ complex multiplications}\approx 64\times 4\text{ real FLOPs}=256\text{ FLOPs}$ (38)
By utilizing this isomorphism, the arithmetic intensity is effectively quartered (neglecting the $O(D)$ overhead of the linear mapping) and the computational workload is shifted from custom sparse kernels to highly-optimized BLAS/LAPACK routines. In the C++ implementation, this transition yields a consistent 25–30% reduction in total forward-pass latency over the bit-masked kernel, allowing Versor to outperform standard Transformers even in wall-clock time.
C.8 Beyond GPUs: Purpose-Built GAPU Architectures
While the bit-masked Triton kernel significantly accelerates GA on existing GPUs, a deeper analysis of the $\mathcal{C}l_{4,1}$ Cayley table reveals that general-purpose hardware is suboptimal for GA workloads, as it allocates silicon area to operations that are structurally redundant in Clifford algebras. This motivates the design of purpose-built Geometric Algebra Processing Units (GAPUs).
C.8.1 Grade-Sparsity Analysis of the Cayley Table
A full geometric product in $\mathcal{C}l_{4,1}$ requires $D^{2}=1024$ multiply-accumulate (MAD) operations. However, most operations in Geometric Product Attention do not require the full product. By analyzing the Cayley table structure from the basis_product_cl41 implementation, the following is observed:
Table 5: Computational cost of structured sub-products in $\mathcal{C}l_{4,1}$, revealing massive grade-sparsity. GPA scoring (scalar product) requires only 3.1% of the full GP’s compute.
Operation Full GP MADs Actual MADs Reduction
Full GP ($A\times B$) 1024 1024 0%
Scalar Product $\langle Q\widetilde{K}\rangle_{0}$ (Scoring) 1024 32 96.9%
Rotor $\times$ Vector (Value Agg.) 1024 80 92.2%
Rotor $\times$ Rotor (RRA Update) 1024 256 75.0%
Sandwich $Rv\widetilde{R}$ (Transform) 2048 336 83.6%
This means the attention scoring phase of GPA—the computational bottleneck—needs only 32 MADs per query-key pair, not 1024. A standard GPU cannot exploit this because its SIMT execution model treats all 32 components uniformly. A purpose-built chip can.
C.8.2 Architecture Simulation Results
Four generations of GAPU architecture are simulated, all transistor-count matched to the NVIDIA A100 ($\sim$54 billion transistors), to ensure a fair silicon-area comparison:
  1. 1.
    GAPU v1 (Parallel Cores): This generation focuses on eliminating the instruction-decode and sign-lookup overhead of general-purpose GPUs. 1,344 independent GA cores are implemented, each containing a fully pipelined Clifford Arithmetic Logic Unit (CALU). The core innovation is the hardcoded sign-logic unit—a combinational circuit of approximately 200 gates—that resolves $Cl(4,1)$ signs in a single clock cycle, effectively zeroing out the latency of Cayley table lookups.
  2. 2.
    CSD (Clifford Systolic Dataflow): Inspired by systolic architectures like the TPU [15], CSD utilizes a $32\times 32$ array of Clifford Processing Elements (CPEs). To minimize data movement, the architecture employs a “weight-stationary” flow: Key and Value multivectors are loaded into a 256MB on-chip Geometric Scratchpad and held resident within the CPE array. Query multivectors pulse through the mesh, where each CPE computes the scalar product scoring via a hardwired FP32 MAC tree. This eliminates 99% of external HBM accesses for standard sequence lengths.
  3. 3.
    GSPA (Grade-Sparse PIM): GSPA represents a shift toward Algebraic Co-Design [21]. By analyzing the grade-sparsity of $Cl(4,1)$, it is observed that attention scoring requires only 32 multiply-adds (3.1% of a full geometric product). GSPA embeds Grade-Aware Clifford Units (GACUs) directly into the HBM3 memory stacks. This Processing-in-Memory (PIM) approach allows scoring to happen in-situ: the result multivector never leaves the individual memory bank, eliding the data movement tax for infinite-context modeling.
  4. 4.
    PHOTON (Wafer-Scale Photonic): PHOTON is a wafer-scale engine (900,000 tiles on a 300mm substrate) utilizing analog silicon photonics [27] for near-instant scoring. Each tile integrates a Mach-Zehnder Interferometer (MZI) array that computes the bilinear scoring operation via optical interference in approximately 10 cycles. The results are fed into a 2GB 3D-stacked SRAM mesh for digital grade-sparse value aggregation, achieving sub-microsecond inference latencies.
Table 6: GAPU Architecture Comparison. Workload: GPA forward pass, Seq=512, Dim=16, Batch=128. All designs are transistor-matched to NVIDIA A100 ($\sim$54B transistors).
Architecture B=1 Latency Speedup (B=128) Energy Eff. Key Innovation
[…]
Figure 4: GAPU architecture evolution: A100 $\to$ Wafer-Scale Photonic Engine. Top-left: absolute latency; top-right: speedup vs A100; bottom-left: energy per forward pass; bottom-right: energy efficiency ratio. All architectures are transistor-matched to ensure fair comparison.
C.8.3 Fabrication and Implementation Regimes
Implementation Costs. The proposed architectures span three distinct fabrication regimes: (1) Commodity CMOS (GAPU v1, CSD): utilizes standard logic fabrication processes with costs estimated at \$5k–10k per unit at volume, potentially offering higher yield than the A100 due to reduced control-logic complexity; (2) Specialized Heterogeneous Integration (GSPA): employs 3D TSV stacking to embed logic within HBM3 layers, with estimated costs of \$20k–30k; and (3) Experimental Wafer-Scale (PHOTON): utilizes full-wafer integration and silicon photonics, placing it in the \$500k+ regime suitable for non-standard deployments.
[…]
Architectural Recommendations. For high-throughput cloud environments and extremely long-context models, GSPA architectures offer a sustainable path by minimizing the “data movement tax” associated with GA multivectors. While CSD exhibits superior raw throughput for standard sequence lengths ($L\leq 4096$) that fit in on-chip SRAM, GSPA serves as the architectural foundation for infinite-context Versor deployments where the cache size exceeds physical scratchpad limits. For low-latency edge applications where sequence length is capped, the CSD architecture provides the optimal balance of throughput and fabrication feasibility. By exploiting the inherent grade structure of $\mathcal{C}l_{4,1}$, these purpose-built processors can effectively circumvent the von Neumann bottlenecks that currently limit Geometric Algebra on general-purpose GPUs.
Reference theoretical models and the full simulation suite are available at GAPU/, within the GitHub repository.
C.9 Proof of Isometric Conformal Embedding
@@
Proposition (Eq. 2): The inner product of two null vectors $X_{1},X_{2}\in\mathcal{C}l_{4,1}$ corresponds to the negative half-squared Euclidean distance between their original points $x_{1},x_{2}\in\mathbb{R}^{3}$.
@@
    Recall the Lifting $\mathcal{K}(\mathbf{x})$: $X=x+\frac{1}{2}x^{2}e_{\infty}+e_{o}$, where $e_{\infty}^{2}=0,e_{o}^{2}=0,$ and $e_{\infty}\cdot e_{o}=-1$.
@@
This verifies that Versor can calculate distances via linear dot products without ever using a $\sqrt{\cdot}$ or nonlinear activation, explaining its efficiency in physics tasks.
C.10 Algorithmic Implementation Specifications
@@ 2 $B_{t}\leftarrow\text{ProjectToBivector}(u_{t},W_{B})$;
C.10.1 Theoretical Roofline Analysis
@@
C.10.2 Arithmetic Intensity ($\mathcal{I}$)
@@
Scenario B: Sparse Table Lookup (Naive GA) Standard GA libraries (Clifford, PyGAE) use a sparse multiplication table.
@@
Scenario C: Versor Bit-Masked Kernel The kernel computes indices/signs in registers.
@@
[…]
@@ D.1 Derivation of Versor Initialization
Problem Formulation: In a standard neural network, weights are initialized to preserve the variance of activations across layers [13]. For a linear layer $y=Wx$, this requires $\text{Var}(y)\approx\text{Var}(x)$. This is generalized to the geometric product $y=Wx$ in $\mathcal{C}l_{4,1}$ (dimension $D=32$). Let $x=\sum_{i=1}^{D}x_{i}e_{i}$ and $W=\sum_{j=1}^{D}w_{j}e_{j}$, where $x_{i},w_{j}$ are independent and identically distributed random variables with mean 0 and variances $\sigma_{x}^{2},\sigma_{w}^{2}$. Recall that $\text{Var}(x)=E[x^{2}]-(E[x])^{2}=E[x^{2}]$. The geometric product is:
@@ D.2 Geometric Product Attention (GPA) Decomposition
    Scalar Part ($\langle\cdot\rangle_{0}$): From Appendix C.9, it is established that $Q\cdot K=-\frac{1}{2}d^{2}$ (approx). Thus, $\text{softmax}(\langle Q\widetilde{K}\rangle_{0})$ recovers the standard RBF-like distance attention used in Vanilla Transformers.
@@
    Torque Modulation: In the current implementation, the bivector magnitude $\|\langle Q\widetilde{K}\rangle_{2}\|$ is incorporated directly into the attention score (see Eq. 4). These properties prioritize interactions not just by proximity (scalar) but by the orientation and “torque” of the plane spanned by the interacting entities. Unlike Euclidean transformers which are blind to relative orientation, GPA attends to the full geometric configuration.
@@ E.3 Gradient Norm Preservation (Backward Stability)
During Backpropagation through Time (BPTT), the gradient of the loss $\mathcal{L}$ with respect to the state $\Psi_{t}$ is computed via the chain rule:
@@
The Jacobian matrix $J_{t}=\frac{\partial\Psi_{t+1}}{\partial\Psi_{t}}$ represents the linear transformation applied to $\Psi_{t}$. In this architecture, this transformation is the left-multiplication by the rotor $\Delta R_{t}$.
@@ E.4 Manifold Adherence of the Cayley Transform
Derivation: Let $B$ be a bivector. In $\mathcal{C}l_{4,1}$, the reversion of a bivector is $\widetilde{B}=-B$. The objective is to prove that $R\widetilde{R}=1$.
@@ Appendix F Experimental Setup and Detailed Results
The library abstracts the complexity of the bit-masked kernels behind a Pythonic API; an example of its use is shown below:
@@ F.1 Chaotic Dynamics (N-Body)
Extended Hamiltonian Details: The Hamiltonian-Versor Hybrid parameterizes the scalar energy function $H(q,p)$ using the Versor backbone. Input state $(q,p)\in\mathbb{R}^{6N}$ is lifted to $\mathcal{C}l_{4,1}$. The output scalar $\langle\Psi_{out}\rangle_{0}$ is treated as the Hamiltonian $\mathcal{H}$. Integration uses a customized 4th-order Runge-Kutta symplectic integrator. This ensures $\frac{d\mathcal{H}}{dt}=0$ by construction (within integrator error).